# SPECIAL RELATIVITY

### v11

@@ 18 Relativity and Electrodynamics
[…]
@@ 18.1 The Dipole Approximation
To address this question, we will analyze the subject more mathematically. For an arbitrary velocity parameter $v/c$, performing covariant calculations of the radiation process is highly challenging. However, in certain cases, significant simplifications arise. One such case is the non-relativistic radiation regime. The non-relativistic asymptotic limit offers a fundamental simplification for covariant calculations. This is because the non-relativistic assumption justifies the dipole approximation, which is of great practical importance. When considering only the dipole component of radiation, all details of the electron trajectory are effectively ignored. Consequently, dipole radiation is completely insensitive to the distinction between covariant and non-covariant particle trajectories  [^59] Similarly, the ultrarelativistic asymptotic limit also simplifies covariant calculations. This is due to the paraxial approximation, which naturally follows from the ultrarelativistic assumption. More details on this effect will be discussed in the next chapter..
[…]
@@
[…]
where $\omega$ is the frequency, $(-e)$ is the negative electron charge and we make use of Gaussian units. A different constant of proportionality in Eq.(LABEL:revwied) and the well-known textbooks is to be ascribed to the use of different units and the definition of the Fourier transform.
@@ 18.2 An Illustrative Example
In contrast, covariant particle tracking—which relies on Lorentz coordinates—yields different results for the electron’s velocity. To explore this, consider a sequence of passive Lorentz transformations used to track the motion of a relativistic electron accelerated by a kicker field. Let $S$ be the lab frame and $S^{\prime}$ a frame comoving with the electron at velocity $\vec{v}$ relative to $S$. Upstream of the kicker, the electron is at rest in the $S^{\prime}$ frame. This implies that $S^{\prime}$ is related to $S$ by a Lorentz boost $L(\vec{v})$, with $\vec{v}$ aligned along the $z$ axis. The boost transforms a four-vector even $X$ in space-time according to $X^{\prime}=L(\vec{v})X$. Now, let us analyze the particle’s dynamics from the perspective of the $S^{\prime}$ frame. In this frame, the electron remains at rest, while the kicker moves toward it with velocity $-\vec{v}$. Due to its motion, the kicker’s magnetic field generates an electric field perpendicular to it. As the kicker interacts with the stationary electron in $S^{\prime}$, the particle experiences a combination of perpendicular electric and magnetic fields.
@@
[…]
@@
According to standard textbooks, in relativistic kinematics, the composition of perpendicular velocities is analyzed using Einstein’s velocity addition theorem. We previously discussed the law of velocity composition in Section 15.5, where it was shown that, within the Lorentz coordinatization, velocity addition is governed by the Wigner rotation. In the ultrarelativistic approximation, a simple result emerges: $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$. This implies that a Lorentz transformation induces a rotation of the particle’s velocity $v_{z}$ by an angle approximately equal to $v_{y}/c$ (Fig. 50). This result highlights that the apparent contradiction is rooted in the presence of the Wigner rotation. As discussed in earlier chapters, the relation $dt^{\prime}=dt/\gamma$ is not valid when addressing the relativistic composition of perpendicular velocities within the Lorentz framework.
It is important to highlight another effective approach to covariant electron tracking. We now revisit the motion of a relativistic electron accelerated by the kicker field, analyzing it from the perspective of an inertial lab observer without changing reference frames. Specifically, we consider an electron in the lab frame undergoing acceleration due to the kicker field—an example of an active boost. In this scenario, we track the evolution of the system over time within a single reference frame.
@@
The simplest synchronization method involves maintaining the same set of uniformly synchronized clocks as in the pre-kick state, without modification. This approach, based on the absolute time (or absolute simultaneity) convention, preserves simultaneity. The standard Galilean velocity addition rule is then applied to track the electron’s motion instant by instant as it moves through a constant magnetic field.
@@
In Chapter 3, we discussed how absolute time coordinatization can be transformed into Lorentz coordinatization. By combining Galilean transformations with variable changes, we ultimately derive the Lorentz transformation within the framework of absolute time coordinatization in the lab frame. To further illustrate this, we now examine how distant lab clocks resynchronize during the electron’s acceleration. This analysis establishes a direct link between the electron’s speed reduction after the kick (in Lorentz coordinates) and time dilation. As previously noted, the Lorentz coordinate system is purely a conceptual construct. Hypothetical clocks are required only for applying Maxwell’s equations in synchrotron radiation calculations.
@@
Suppose that, upstream of the kicker, we choose a Lorentz coordinate system in the lab frame. Immediately after the electron enters the magnetic field, its velocity changes by an infinitesimal amount $d\vec{v}$ along the $y$-axis. At this initial stage, Eq. (22) allows us to express the differential $d\vec{v}$ in terms of the differential $dt$ within the Lorentz coordinate system defined upstream of the kicker. If clock synchronization is fixed, this corresponds to adopting the absolute time convention. To maintain Lorentz coordinates in the lab frame—as discussed previously—we must perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case arises when the kick angle $\theta_{k}$ is very small, allowing us to work up to second-order terms in $(\theta_{k}\gamma)^{2}$. This restriction greatly simplifies the calculations for two reasons. First, relativistic corrections due to the composition of non-collinear velocity increments only appear at order $(\gamma\theta_{k})^{3}$ and can therefore be neglected. Second, time dilation effects also enter only at higher orders. Thus, Eq. (22) enables us to express the small velocity change $\Delta\vec{v}$ due to the kick in the initial Lorentz coordinate system and defer clock resynchronization until after the kicker. Consequently, the post-kick motion can be described as a composition of two Lorentz boosts in the perpendicular $y$ and $z$ directions. The first boost imparts a velocity $v\theta_{k}\vec{e}_{y}$ along the $y$-axis, while the second applies an additional velocity $-(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis. The second-order approximation ensures that the two boosts commute.
@@
This analysis is performed in the limit of a small kick angle $\gamma v_{y}/c\ll 1$. Nevertheless, even in this simplified case and for a single electron, we are able to clearly demonstrate the distinction between covariant and non-covariant particle trajectories.
@@ 18.4 Redshift of the Synchrotron Radiation Critical Frequency
It is worth noting, however, that there exists an alternative and equally satisfactory explanation for the redshift. This interpretation employs a non-covariant framework, which considers non-covariant particle trajectories and applies Galilean transformations to the electromagnetic field equations. In this approach, the electron velocity remains unchanged according to non-covariant particle tracking. However, Maxwell’s equations are not invariant under Galilean transformations, leading to a change in the speed of light. Specifically, the velocity of light increases from $c$ (before the kick) to $c(1+\theta_{k}^{2}/2)$ (after the kick). This variation arises because, under the absolute time convention, clocks are not resynchronized after the kick. As a result, the speed of light differs from the electrodynamical constant $c$. Nevertheless, the ratio of the electron velocity to the speed of light remains invariant across different synchronization conventions—that is, it does not depend on the method of distant clock synchronization or the rate of the clocks. Our calculations demonstrate that both covariant and non-covariant treatments—when the correct coupling between fields and particles is used—yield the same prediction for the redshift. This prediction is convention-invariant and depends solely on the dimensionless parameter $v_{coord}/c_{coord}=v(1-\theta^{2}_{k}/2)/c=v/[c(1+\theta^{2}_{k}/2)]$, where $v_{coord}$ is the coordinate velocity of the electron, and $c_{coord}$ is the coordinate velocity of light.
@@
18.5 Discussion
We now wish to emphasize that when an electron is accelerated in the lab frame by a kicker, this acceleration is encoded in its covariant trajectory. A compelling question arises: Where, physically, is the information about the electron’s acceleration located when using absolute time coordinatization?
[…]
This apparent paradox—of an unchanged velocity despite acceleration—can be explained through dynamical reasoning involving the electron’s non-covariant trajectory. Without entering into a full derivation, we summarize the key insight as follows:
For a rapidly moving electron, the transverse electric and magnetic fields are nearly equal in magnitude and orthogonal to each other—indistinguishable from the fields of an actual radiation beam. On a microscopic level, when the electron passes through a kicker, its electromagnetic fields are perturbed and now carry imprints of that acceleration.
Under the framework of classical (Galilean) kinematics, the orientation of the virtual radiation phase front remains unchanged. However, since Maxwell’s equations are not invariant under Galilean transformations—as discussed throughout this article—the adoption of such kinematics necessitates the use of anisotropic field equations. Consequently, although the phase front remains planar, the direction of propagation is no longer perpendicular to it. In other words, the radiation beam’s direction of motion diverges from the normal to its phase front.
Within the absolute time synchronization convention, electrodynamics predicts that the virtual radiation beam propagates along the direction of the kick, with a phase front tilt $\theta_{k}$. This tilt is the crucial clue to resolving the paradox. The key insight is this: The information about the electron’s acceleration is not lost—it is embedded in the perturbation of the electron’s self-fields. Mathematically, this information is encoded in the tilt of the phase front of the virtual radiation beam.
Let us now briefly consider some additional, intriguing cases. Previously, we examined a simple scenario in which the kick angle is small. In such cases, when an electron passes through a weak kicker, it emits no appreciable synchrotron radiation. More precisely, the radiation is indistinguishable from the electron’s intrinsic electromagnetic fields. We can therefore assume that the electron remains in a “field-dressed” state — still accompanied by its usual virtual radiation fields — after passing through the weak kicker.
Now, let us turn to a different situation. Suppose an ultrarelativistic electron in the laboratory frame is deflected by a dipole magnetic field with a kick angle much larger than $1/\gamma$, This setup corresponds to a typical synchrotron radiation scenario.
At the exit of the strong kicker, the electron emerges effectively “naked” or “field-free,” meaning it is no longer accompanied by its usual virtual field structure. A process then begins in which the “field-dressed” state is gradually re-established — that is, the characteristic electromagnetic fields associated with a fast-moving charge begin to reform — over a certain formation length in the downstream straight section.
At this point, information about the electron’s recent acceleration is carried away by the emitted synchrotron radiation. However, the electron’s original self-fields and its acceleration history are effectively erased in the strong kick. As a result, the electromagnetic fields surrounding the electron downstream of the kicker no longer encode this history. In other words, the electron’s covariant trajectory beyond the kicker contains no imprint of the preceding acceleration. We will revisit this topic in more detail in the next chapter.
@@ 19.2 Paraxial Approximation of the Radiation Field
[…]
@@
We will now replace all vectors by their components to obtain directional dependency of the synchrotron radiation. The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis. Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the $x-z$ plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the $y-z$ plane. The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$. We consider the motion in a static magnetic field. According to conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path. The transverse components of the envelope of the field in Eq. (LABEL:rrevwied) in the far zone and paraxial approximation finally becomes
[…]
@@ 19.3.1 Conventional Theory
[…]
@@
[…]

### v12

@@ 19 Relativity and Electrodynamics
In this book, we emphasize the significant role of the Wigner rotation as a regulator of the velocity addition law. As previously discussed, the four-velocity of an accelerating particle moving along a curved path in the Lorentz lab frame cannot, in general, be decomposed as $u=(c\gamma,\vec{v}\gamma)$.  [^61] Consider, for instance, a typical textbook treatment [72], which discusses the projection of an arbitrary world line onto the Lorentz lab frame basis: ”The particle, which is assumed to curry the charge $e$, creates the current density $j(x)=ec\int d\tau u(y)\delta^{4}(y-x(\tau))$. […] Furthermore, in any frame of reference $K$, one recovers the expected expressions for the charge and current densities by integrating over $\tau$ by means of relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time and using the formula $\delta(y_{0}-x_{0}(\tau))=\delta(ct-ct^{\prime})=\delta(t-t^{\prime})/c$, $j_{0}(t,y)=ce\delta^{(3)}(y-x(t))\equiv c\rho(t,y)$, $j^{i}(t,y)=ev^{i}(t)\delta^{(3)}(y-x(t))$, i=1,2,3i=1,2,3”. The presentation of the time component simply as the relation $d\tau=dt^{\prime}/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
@@ 19.1 The Dipole Approximation
To address this question, we will analyze the subject more mathematically. For an arbitrary velocity parameter $v/c$, performing covariant calculations of the radiation process is highly challenging. However, in certain cases, significant simplifications arise. One such case is the non-relativistic radiation regime. The non-relativistic asymptotic limit offers a fundamental simplification for covariant calculations. This is because the non-relativistic assumption justifies the dipole approximation, which is of great practical importance. When considering only the dipole component of radiation, all details of the electron trajectory are effectively ignored. Consequently, dipole radiation is completely insensitive to the distinction between covariant and non-covariant particle trajectories.  [^62] Similarly, the ultrarelativistic asymptotic limit also simplifies covariant calculations. This is due to the paraxial approximation, which naturally follows from the ultrarelativistic assumption. More details on this effect will be discussed in the next chapter.
We want now to solve electrodynamics equations mathematically in a general way and consider the radiation associated with the succeeding terms in (multi-pole) expansion of the field in powers of the ratio $v/c$. Radiation theory is naturally developed in the space-frequency domain, as one is usually interested in radiation properties at a given position in space and at a certain frequency. [^63] In this book we define the relation between temporal and frequency domain via the following definition of Fourier transform pair: $\bar{f}(\omega)=\int_{-\infty}^{\infty}dt~f(t)\exp(i\omega t)\leftrightarrow f(t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}d\omega\bar{f}(\omega)\exp(-i\omega t)$.
@@
$\vec{\bar{E}}(\vec{r}_{0},\omega)$ $=$ ${i\omega e\over{cr_{0}}}\exp\left[-\frac{i\omega}{c}\vec{n}\cdot\vec{r}_{0}\right]\int_{-\infty}^{\infty}dt~{\vec{n}\times\left[\vec{n}\times{\vec{\beta}(t)}\right]}\exp\left[i\omega\left(t+\frac{\vec{n}\cdot\vec{r}(t)}{c}\right)\right]$ (36)
where $\omega$ is the frequency, (e)(-e) is the negative electron charge and we make use of Gaussian units. A different constant of proportionality in Eq.(LABEL:revwied) and the well-known textbooks is to be ascribed to the use of different units and the definition of the Fourier transform.
@@ 19.2 An Illustrative Example
In contrast, covariant particle tracking—which relies on Lorentz coordinates—yields different results for the electron’s velocity. To explore this, consider a sequence of passive Lorentz transformations used to track the motion of a relativistic electron accelerated by a kicker field. Let $S$ be the lab frame and $S^{\prime}$ a frame comoving with the electron at velocity $\vec{v}$ relative to $S$. Upstream of the kicker, the electron is at rest in the $S^{\prime}$ frame. This implies that $S^{\prime}$ is related to $S$ by a Lorentz boost $L(\vec{v})$, with $\vec{v}$ aligned along the $z$ axis. The boost transforms a four-vector event $X$ in space-time according to $X^{\prime}=L(\vec{v})X$. Now, let us analyze the particle’s dynamics from the perspective of the $S^{\prime}$ frame. In this frame, the electron remains at rest, while the kicker moves toward it with velocity $\vec{v}$. Due to its motion, the kicker’s magnetic field generates an electric field perpendicular to it. As the kicker interacts with the stationary electron in $S^{\prime}$, the particle experiences a combination of perpendicular electric and magnetic fields.
@@
We begin by analyzing the textbook treatment of the composition of motions. Let SS" be a reference frame fixed with respect to a particle located downstream of the kicker. As is well known, non-collinear Lorentz boosts do not commute. However, in our second-order approximation, we can neglect the difference between $\gamma v_{y}/c$ and $\gamma_{z}v_{y}/c$, where $\gamma_{z}=1/\sqrt{1-v_{z}^{2}/c^{2}}$. Here $v_{z}=v(1-\theta_{k}^{2}/2)$ and $\theta_{k}=v_{y}/v=v_{y}/c$ in the ultrarelativistic regime of interest. Therefore, we can approximate the transformation from coordinates $X^{\prime}$ in frame $S^{\prime}$ to coordinates XX" in SS" using a sequence of two commuting non-collinear Lorentz boosts: $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})X^{\prime}=L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e^{\prime}}_{y}v^{\prime}_{y})X^{\prime}$, where $\vec{e^{\prime}}_{y}$ and $\vec{e^{\prime}}_{z}$ are unit vectors along the $x^{\prime}$ and $z^{\prime}$ axis respectively. It is important to note that, as seen by an observer in $S^{\prime}$, the axes of frameSS" remain parallel to those of $S^{\prime}$. The full transformation from the lab frame $S$ to the downstream particle frame SS" can thus be written as: $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)X$. In the special case where the velocities are collinear, the composition of boosts simplifies: $L(\vec{e^{\prime}}_{z}v^{\prime}_{z})L(\vec{e}_{z}v)=L(\vec{e}_{z}v_{z})$, leading to the overall transformation: $X"=L(\vec{e^{\prime}}_{y}v^{\prime}_{y})L(\vec{e}_{z}v_{z})X$. Textbooks typically emphasize that, within the Lorentz framework, a magnetic field can only alter the direction of an electron’s motion, but not its speed.
@@
According to standard textbooks, in relativistic kinematics, the composition of perpendicular velocities is analyzed using Einstein’s velocity addition theorem. We previously discussed the law of velocity composition in Section 16.6, where it was shown that, within the Lorentz coordinatization, velocity addition is governed by the Wigner rotation. In the ultrarelativistic approximation, a simple result emerges: $v_{tot}=v_{z}=v[1-v_{y}^{2}/(2c^{2})]$. This implies that a Lorentz transformation induces a rotation of the particle’s velocity $v_{z}$ by an angle approximately equal to $v_{y}/c$ (Fig. 50). This result highlights that the apparent contradiction is rooted in the presence of the Wigner rotation.
@@
Now let us revisit the kinematics of a relativistic electron accelerated by the kicker field, analyzing it from the perspective of an inertial lab observer without changing reference frames. Specifically, we consider an electron in the lab frame undergoing acceleration due to the kicker field. The simplest synchronization method involves maintaining the same set of uniformly synchronized clocks as in the pre-kick state, without modification. This approach, based on the absolute time (or absolute simultaneity) convention, preserves simultaneity. The standard Galilean velocity addition rule is then applied to track the electron’s motion instant by instant as it moves through a constant magnetic field.
@@
By combining Galilean transformations with variable changes, we ultimately derive the Lorentz transformation within the framework of absolute time coordinatization in the lab frame. This analysis establishes a direct link between the electron’s speed reduction after the kick (in Lorentz coordinates) and time dilation.
@@
Suppose that, upstream of the kicker, we choose a Lorentz coordinate system in the lab frame. Immediately after the electron enters the magnetic field, its velocity changes by an infinitesimal amount $d\vec{v}$ along the $y$-axis. At this initial stage, Eq. (22) allows us to express the differential $d\vec{v}$ in terms of the differential $dt$ within the Lorentz coordinate system defined upstream of the kicker. If clock synchronization is fixed, this corresponds to adopting the absolute time convention. To maintain Lorentz coordinates in the lab frame—as discussed previously—we must perform a clock resynchronization by introducing an infinitesimal time shift. The simplest case arises when the kick angle $\theta_{k}$ is very small, allowing us to work up to second-order terms in $(\theta_{k}\gamma)^{2}$. This restriction greatly simplifies the calculations for two reasons. First, relativistic corrections due to the composition of non-collinear velocity increments only appear at order $(\gamma\theta_{k})^{3}$ and can therefore be neglected. Second, time dilation effects also enter only at higher orders. Thus, Eq. (22) enables us to express the small velocity change $\Delta\vec{v}$ due to the kick in the initial Lorentz coordinate system and defer clock resynchronization until after the kicker. Consequently, the post-kick motion can be described as a composition of two Lorentz boosts in the perpendicular $y$ and $z$ directions. The first boost imparts a velocity $v\theta_{k}\vec{e}_{y}$ along the $y$-axis, while the second applies an additional velocity $(v\theta_{k}^{2}/2)\vec{e}_{z}$ along the $z$ axis. The second-order approximation ensures that the two boosts commute.
@@ 20.2 Paraxial Approximation of the Radiation Field
$\vec{\bar{E}}(\vec{r}_{0},\omega)$ $=$ $-{i\omega e\over{cr_{0}}}\exp\left[-\frac{i\omega}{c}\vec{n}\cdot\vec{r}_{0}\right]\int_{-\infty}^{\infty}dt~[\vec{\beta}(t)-\vec{n}]\exp\left[i\omega\left(t+\frac{\vec{n}\cdot\vec{r}(t)}{c}\right)\right]$ (38)
@@
We will now replace all vectors by their components to obtain directional dependency of the synchrotron radiation. The emission angle $\theta=\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ is taking with respect to the $z$-axis. Here $\theta_{x}=x_{0}/z_{0}$ is the observation angle projected onto the xzx-z plane, $\theta_{y}=y_{0}/z_{0}$ is the observation angle projected onto the yzy-z plane. The components of the unit vector $\vec{n}$ can be approximated by $n_{z}=1-\theta_{x}^{2}/2-\theta_{y}^{2}/2$, $n_{x}=\theta_{x}$, $n_{y}=\theta_{y}$, so $\vec{n}\cdot(\vec{r}_{0}-\vec{r})=(z_{0}-z)(1-\theta_{x}^{2}/2-\theta_{y}^{2}/2)-x\theta_{x}-y\theta_{y}$. We consider the motion in a static magnetic field. According to conventional particle tracking the magnitude of the velocity is constant and is equal $v=ds/dt=\mathrm{const.}$, where $s(z)$ is the longitudinal coordinate along the path. The transverse components of the envelope of the field in Eq. (LABEL:rrevwied) in the far zone and paraxial approximation finally becomes
$\vec{\widetilde{{E}}}(z_{0},\vec{r}_{0},\omega)$ $=$ $-{i\omega e\over{c^{2}}z_{0}}\int_{-\infty}^{\infty}dz^{\prime}{\exp{\left[i\Phi_{T}\right]}}\left[\left({v_{x}(z^{\prime})\over{c}}-\theta_{x}\right){\vec{e_{x}}}+\left({v_{y}(z^{\prime})\over{c}}-\theta_{y}\right){\vec{e_{y}}}\right]$ (43)
@@ 20.3.1 Conventional Theory
${\vec{\widetilde{E}}}={i\omega e\over{c^{2}z_{0}}}\int_{-L/2}^{L/2}dz^{\prime}{\exp\left[i\Phi_{T}\right]\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]}\left[{K\over{\gamma}}\sin\left(k_{w}z^{\prime}\right)\vec{e}_{x}+\vec{\theta}\right]~.$ (50)
@@
${\vec{\widetilde{E}}}=-\frac{K\omega e}{2c^{2}z_{0}\gamma}A_{JJ}\exp\left[i\frac{\omega\theta^{2}z_{0}}{2c}\right]\int_{-L/2}^{L/2}dz^{\prime}\exp\left[i\left(C+{\omega\theta^{2}\over{2c}}\right)z^{\prime}\right]\vec{e}_{x}~,$ (55)